OpenStudy (anonymous):

proof: (1+tanx) / (1+cotx) = tan x

OpenStudy (lgbasallote):

tan x = sin/cos cot = cos/sin does that help?

OpenStudy (mertsj):

\[\frac{1+\tan x}{1+\frac{1}{\tan x}}=\frac{\tan x+\tan ^2x}{\tan x+1}=\frac{\tan x(1+\tan x)}{\tan x+1}=\tan x\]

OpenStudy (anonymous):

1+tanx/1+cotx=tanx: multiply the denominator by both sides: 1+tanx=tanx(1)+tanx(cotx) tanx(1)=tanx and tanx(cotx) equals one so it turns out to be: tanx +1=tanx + 1 thats it:)

OpenStudy (mertsj):

@xXDWEAHXx87 Can't work across the equal sign when proving an identity. The normal rules of equations do not apply.

OpenStudy (anonymous):

o sry

OpenStudy (mertsj):

no problem

OpenStudy (anonymous):

but it still works though

OpenStudy (mertsj):

yep

OpenStudy (mertsj):

It works to cheat at poker too

OpenStudy (anonymous):

but ur way is correct and more awesome looking so good job